OpenStudy (anonymous):

If sin θ = 2/x in the first quadrant, what is the expression for the hypotenuse of the right triangle? What about cos θ, sec θ, and cot θ?

OpenStudy (anonymous):

Sin = opp/hyp Right? So what's your hypotenuse here?  Sorry meant 2 not 12. Can you solve the rest?

OpenStudy (anonymous):

sec= 1/cos and cot=1/tan

OpenStudy (ajprincess):